# Orak: A Foundational Benchmark for Training and Evaluating LLM Agents on Diverse Video Games

### v2

@@ Abstract
Large Language Model (LLM) agents are reshaping the game industry, by enabling more intelligent and human-preferable characters. Yet, current game benchmarks fall short of practical needs: they lack evaluations of diverse LLM capabilities across various game genres, studies of agentic modules crucial for complex gameplay, and fine-tuning datasets to adapt pre-trained LLMs into gaming agents. To fill these gaps, we present Orak, a benchmark for training and evaluating LLM agents across 12 popular video games spanning all major genres. Using a plug-and-play interface built on Model Context Protocol (MCP), Orak supports systematic and reproducible studies of agentic modules in varied game scenarios. We further release a fine-tuning dataset of expert LLM gameplay trajectories covering multiple genres, turning general LLMs into effective game agents. Orak offers a united evaluation framework, including game leaderboards, LLM battle arenas, and in-depth analyses of input modality, agentic strategies, and fine-tuning effects, establishing a foundation towards versatile gaming agents. Code is available at https://github.com/krafton-ai/Orak.
@@ 1 Introduction
Large Language Model (LLM) agents are revolutionizing various industries (Wang et al., 2024a), and well-established benchmarks play a key role in unlocking their abilities on complex tasks, e.g., Coding (Hendrycks et al., 2021, Zhuo et al., 2024), Web Search (Liu et al., 2023, Pan et al., 2024, Levy et al., 2024), and Scientific Research (Mühlbacher et al., 2024, Zhang et al., 2025). In game industry, there is growing interest in using LLM agents to enhance user game experiences e.g., more intelligent non-player characters (NPCs), monsters, or companions (NVIDIA, 2025). In response, many benchmarks have been proposed to assess the ability of LLMs to play games (Hu et al., 2024b).
@@
Our benchmark offers comprehensive evaluation dimensions, including game score leaderboards, competitive LLM battle arenas, and in-depth analyses of visual input state, agentic strategies, and fine-tuning effects. Extensive experiments on Orak with 15 LLMs reveal that (1) proprietary LLMs achieve superior performance across games with significant gaps from open-source LLMs, (2) their performance gap becomes narrow in battle scenarios, (3) proprietary LLMs benefit from extended agentic workflows, while open-source LLMs show limited gains, (4) visual states often hinder gameplay performance yet, and (5) fine-tuning enables effective transfer of gameplay meta-knowledge from larger LLMs to smaller ones, leading to generalization in intra-game, out-of-distribution (OOD) game, and non-game unseen scenarios like math and web interaction. We believe that Orak not only establishes a foundation for building gaming agents but also serves as a critical benchmark for evaluating general LLMs on realistic, long-horizon decision-making tasks.
@@ 2 Related Work
As gameplay requires complex cognitive abilities, e.g., context understanding, logical reasoning, and error handling, several recent benchmarks have sought to evaluate LLMs or Vision Language Models (VLMs) on games (Hu et al., 2024b). GAMA-Bench (Huang et al., 2024), GameBench (Costarelli et al., 2024), GameArena (Hu et al., 2024a), and SmartPlay (Wu et al., 2023) focus on text-based games, assessing LLMs’ ability to navigate and reason on textual environments. Barlog (Paglieri et al., 2024) and LVLM-Playground (Wang et al., 2025) are mainly based on 2D-grid games, such as TicTacToe and Chess, to evaluate the spatial and visual reasoning capabilities of LLMs/VLMs. Using video games, Cradle (Tan et al., 2024) evaluates VLMs in 1 adventure and 3 simulation games, V-MAGE (Zheng et al., 2025) assesses VLMs on 5 action games, DSGBench (Tang et al., 2025) validates LLMs on 6 strategic games, and LMGame-Bench (Hu et al., 2025) mainly focus on puzzle games. Despite their contributions, existing benchmarks lack full coverage of game genres, omit in-depth studies on agentic modules, often rely on visual inputs, and under-explore the way to align pre-trained LLMs into versatile game agents. Table 1 summarizes the key characteristics of game benchmarks.
@@
| LVLM-Playground (Wang et al., 2025) | 2D-grid | ✗ | 6 | VLM | ✗ | ✗ |
@@ 4 Fine-tuning: Aligning Pre-trained LLMs into Game Agents
| Data example from Reflection module | Data example from Action module | ||||||||||||
@@ 5.4 Ablation Study for Agentic Modules
Table 4 shows the ablation results of LLaMA-3.2-3B and GPT-4o across 4 agent strategies. Interestingly, the impact of adding agentic modules to gameplay performance differs between the two LLMs. For GPT-4o, the inclusion of agentic modules consistently improves gameplay performance; ‘reflection-planning’ agent achieves the best average ranking of 2.2, followed by ‘reflection’ and ‘planning’ agents with rankings of 3.1 and 3.3, and ‘zero-shot’ agent with the lowest ranking of 3.4. However, LLaMA-3.2-3B does not follow this trend. The ‘reflection’ agent shows the highest average ranking of 5.6, while ‘reflection-planning’ follows with 6.1, although it adds the planning module. This indicates that, for relatively smaller LLMs like LLaMA-3.2-3B, adding agentic modules may increase the complexity of the prompt, hindering their decision-making accuracy. These results suggest that the optimal agentic strategy may depend on the inherent capability of the LLM.
@@ 5.6 Effect of Fine-tuning
To study the effect of fine-tuning, we consider three types of generalization: Intra-game, OOD-game, and Non-game. Intra-game generalization evaluates whether an LLM can adapt to unseen scenarios within the same game, e.g., new stages or characters. OOD-game generalization evaluates whether fine-tuning on a specific set of games enables the model to act better on unseen games. Non-game generalization evaluates whether our fine-tuning set with agent gameplay trajectory helps the model to perform better on non-game tasks like math or web navigation. For Intra-game and Non-game generalizations, we fine-tune LLMs on all 12 games, while for OOD-game generalization, we split 12 games in to 10 training games and 2 test games, i.e., Super Mario and 2048. See Appendix P for detailed fine-tuning configurations and unseen scenario details in intra-game generalization.
@@
As shown in Table 7, fine-tuned Llama-3.2-1B/3B show better generalization to unseen scenarios than the pretrained ones; in 3 out of 5 games, both fine-tuned models outperform their pretrained counterparts. The performance gain largely comes from the model learning to generate valid actions more reliably after fine-tuning, especially in environments where the pretrained model frequently failed to act meaningfully. However, this approach appears insufficient to significantly enhance the spatial reasoning abilities of smaller models. This is evident in Baba Is You, where fine-tuned models struggle to construct the winning condition required to score above 20.0.
@@ Conclusion.
We note that Orak provides environments favorable for LLM reasoning and decision-making by pre-processing game states into structured text where information unnecessary to gameplay is hidden. This may offer insights for providers seeking to deploy resource-efficient LLMs on games. We leave a complementary direction, providing the full in-game states with rich uncurated texts to LLMs, for future work. Further discussions, including cost and licensing, are provided in Appendix R.
@@ C.1 Game Description for Street Fighter III
Observation-to-Text Conversion. The Diambra environment offers a convenient interface for extracting the game state from Street Fighter III. Through this interface, we obtain the latest game frame at a resolution of 224×384, along with key state information such as remaining time, player and opponent health, super bar gauge, super count, stun bar gauge, and stun status. However, a critical aspect in fighting games is understanding the relative positions of the characters, which is not directly provided by the Diambra environment. To address this limitation, we employ a lightweight YOLOv11 object detection model (Khanam & Hussain, 2024) to extract the relative positions of the two characters from the game frame. Based on the computed distance, we classify the spatial relationship into three discrete categories—very close, close, and far—and incorporate this information into the agent’s user prompt.
@@ C.3 Evaluation Metric for Street Fighter III
We adopt a Bradley-Terry model formulation (Bradley & Terry, 1952) for Elo estimation, where each model’s rating is iteratively optimized using gradient ascent on the log-likelihood of observed outcomes. The gradient is computed based on the expected win probabilities derived from current ratings, using the standard Elo transformation:
@@ C.5 Result for Street Fighter III
All reported results in Tables 11–12 are computed as the mean and standard deviation over three to five independent runs, using a ‘zero-shot’ agent for each model configuration. To establish a baseline for comparison, we additionally included a random agent that selects an arbitrary action uniformly at random. This agent was evaluated over 30 episodes, and its performance is summarized in tables.
@@ I.5 Result for Minecraft
Using these difference-based scores, we then compute Elo ratings for all agents following the Bradley-Terry model (Bradley & Terry, 1952), following the approach used in StreetFighter III multi-agent evaluation in Section C.3.
@@
All reported results in Tables 45–45 are computed as the mean and standard deviation over five independent runs, using a ‘zero-shot’ agent for each model configuration. To establish a baseline for comparison, we additionally included a random agent that selects one of the four possible actions (up, down, left, right) uniformly at random. This agent was evaluated over 50 episodes, and its performance is summarized in Tables.

### v3

@@ Abstract
Large Language Model (LLM) agents are reshaping the game industry, by enabling more intelligent and human-preferable characters. Yet, current game benchmarks fall short of practical needs: they lack evaluations of diverse LLM capabilities across various game genres, studies of agentic modules crucial for complex gameplay, and fine-tuning datasets to adapt pre-trained LLMs into gaming agents. To fill these gaps, we present Orak, a benchmark for training and evaluating LLM agents across 12 popular video games spanning all major genres. Using a plug-and-play interface built on Model Context Protocol (MCP), Orak supports systematic and reproducible studies of agentic modules in varied game scenarios. We further release a fine-tuning dataset of expert LLM gameplay trajectories covering multiple genres, turning general LLMs into effective game agents. Orak offers a united evaluation framework, including game leaderboards, LLM battle arenas, and ablation studies of input modality, agentic strategies, and fine-tuning effects, establishing a foundation towards versatile gaming agents. Code and datasets are available at https://github.com/krafton-ai/Orak and https://huggingface.co/datasets/KRAFTON/Orak.
@@ 1 Introduction
Large Language Model (LLM) agents are revolutionizing various industries (Wang et al., 2024a), and well-established benchmarks play a key role in unlocking their abilities on complex tasks, e.g., Coding (Hendrycks et al., 2021, Zhuo et al., 2024), Web Search (Liu et al., 2023, Pan et al., 2024), and Scientific Research (Mühlbacher et al., 2024, Zhang et al., 2025). In the game industry, there is growing interest in using LLM agents to enhance user game experiences, e.g., more intelligent non-player characters (NPCs), monsters, or companions (NVIDIA, 2025). Also, games that simulate dynamic and uncertain real-world scenarios can serve as effective test-beds for evaluating agents’ high-level decision-making, i.e., system 2 reasoning (Li et al., 2025). In response, many benchmarks have been proposed to assess the ability of LLMs by playing games (Hu et al., 2024b, Ahn et al., 2025).
@@
Our benchmark offers comprehensive evaluation dimensions, including game score leaderboards, competitive LLM battle arenas, and ablation studies of visual input state, agentic strategies, and fine-tuning effects. Extensive experiments on Orak with 15 LLMs reveal that (1) proprietary LLMs achieve superior performance across games with significant gaps from open-source LLMs, (2) their performance gap becomes narrow in battle scenarios, (3) proprietary LLMs benefit from extended agentic workflows, while open-source LLMs show limited gains, (4) models fail to extract sufficient value from visual inputs yet, and (5) fine-tuning enables effective transfer of gameplay meta-knowledge from larger LLMs to smaller ones, leading to generalization in intra-game, out-of-distribution (OOD) game, and non-game unseen scenarios like math and web interaction. We believe that Orak not only establishes a foundation for building gaming agents but also serves as a critical benchmark for evaluating general LLMs on realistic, long-horizon decision-making tasks.
@@ 2 Related Work
As gameplay requires complex cognitive abilities, e.g., context understanding, logical reasoning, and error handling, several recent benchmarks have sought to evaluate LLMs or Vision Language Models (VLMs) on games (Hu et al., 2024b). GAMA-Bench (Huang et al., 2024), GameBench (Costarelli et al., 2024), GameArena (Hu et al., 2024a), and SmartPlay (Wu et al., 2023) focus on text-based games, assessing LLMs’ ability to navigate and reason on textual environments. Barlog (Paglieri et al., 2024) and LVLM-Playground (Wang et al., 2025a) are mainly based on 2D-grid games, such as TicTacToe and Chess, to evaluate the spatial and visual reasoning capabilities of LLMs/VLMs. Using video games, Cradle (Tan et al., 2024) evaluates VLMs in 1 adventure and 3 simulation games, V-MAGE (Zheng et al., 2025) assesses VLMs on 5 action games, DSGBench (Tang et al., 2025) validates LLMs on 6 strategic games, and LMGame-Bench (Hu et al., 2025) mainly focus on puzzle games. Despite their contributions, existing benchmarks lack full coverage of game genres, omit in-depth studies on agentic modules, often rely on visual inputs, and under-explore the way to align pre-trained LLMs into versatile game agents. Table 1 summarizes the key characteristics of game benchmarks.
@@
| LVLM-Playground (Wang et al., 2025a) | 2D-grid | ✗ | 6 | VLM | ✗ | ✗ |
@@ 4 Fine-tuning: Aligning Pre-trained LLMs into Game Agents
| Data example from Reflection module | Data example from Action module | ||
| Analyze Mario’s past action using state difference and provide critiques for improving his action. You should only respond in the format as below: Self-reflection (Describe self-reflection here) | | (Retrieve the self-reflection from memory) Analyze the current game state, and decide the best action. You should only respond in the format as below: Action Jump Level: n (where n is an integer from 0 to 6) | |
| | Past Game State Mario at (100,100), Bricks at (120, 100), (120, 150) Current Game State Mario at (100,100), Bricks at (120, 100), (120, 150) | | Game State Mario at (100,100) Position of all objects: - Bricks at (120, 100), (120, 150) |
| | Self-reflection Mario is blocked by bricks. Jump higher to get past. | | Action Jump Level: 6. |
@@ 5.4 Ablation Study for Agentic Modules
Table 4 shows the ablation results of Llama-3.2-3B and GPT-4o across 4 agent strategies. Interestingly, the impact of adding agentic modules to gameplay performance differs between the two LLMs. For GPT-4o, the inclusion of agentic modules consistently improves gameplay performance; ‘reflection-planning’ agent achieves the best average ranking of 2.2, followed by ‘reflection’ and ‘planning’ agents with rankings of 3.1 and 3.3, and ‘zero-shot’ agent with the lowest ranking of 3.4. However, Llama-3.2-3B does not follow this trend. The ‘reflection’ agent shows the highest average ranking of 5.6, while ‘reflection-planning’ follows with 6.1, although it adds the planning module. This indicates that, for relatively smaller LLMs like Llama-3.2-3B, adding agentic modules may increase the complexity of the prompt, hindering their decision-making accuracy. These results suggest that the optimal agentic strategy may depend on the inherent capability of the LLM. See Appendices C-N for more detailed ablation analysis for each game.
@@ 5.6 Effect of Fine-tuning
To study the effect of fine-tuning, we consider three types of generalization: Intra-game, OOD-game, and Non-game. Intra-game generalization evaluates whether an LLM can adapt to unseen scenarios within the same game, e.g., new stages or characters. OOD-game generalization evaluates whether fine-tuning on a specific set of games enables the model to act better on unseen games. Non-game generalization evaluates whether our gameplay fine-tuning set helps the model to perform better on non-game tasks like math or web navigation. For Intra-game and Non-game generalizations, we fine-tune LLMs on all 12 games, while for OOD-game generalization, we split 12 games in to 10 training games and 2 test games, i.e., Super Mario and 2048. Additionally, to study the effect of data quality and scaling, we perform fine-tuning on high-score (original), low-score (samples with the lowest game score), and their mixed datasets (doubling the data volume). See Appendix P for detailed fine-tuning configurations and unseen scenario details in intra-game generalization.
@@
As shown in Table 7, fine-tuned Llama-3.2-1B/3B show better generalization to unseen scenarios than the pretrained ones; in 3 out of 5 games, both fine-tuned models outperform their pretrained counterparts. The performance gain largely comes from the model learning to generate valid actions more reliably after fine-tuning, especially in environments where the pretrained model frequently failed to act meaningfully. However, this approach appears insufficient to significantly enhance the spatial reasoning abilities of smaller models. This is evident in Baba Is You, where fine-tuned models struggle to construct the winning condition required to score above 20.0. Also, since StarCraft II requires complex strategic reasoning to defeat enemies, the strategic patterns learned from the fine-tuning set were not sufficient for the model to win on new maps.
@@
Effect of Data Quality and Scaling.
Table 8(a) shows the effect of data quality and scaling in fine-tuning generalization. Overall, high-score dataset alone is the most effective, while the low-score dataset contributes little or even degrades performance in certain cases (e.g., Darkest Dungeon). Simply scaling up the dataset by mixing high- and low-score samples does not yield a clear improvement, indicating that data quality, rather than sheer quantity, plays a more critical role in achieving better generalization.
| Data | SF3 | DarkestD | StarCraft2 | SlaySpire | BabaIsYou |
| No fine-tuning | 12.011.0 | 87.29.5 | 0.00.0 | 0.00.0 | 20.00.0 |
| high-score (110K) | 40.07.1 | 92.00.2 | 0.00.0 | 10.71.2 | 20.00.0 |
| low-score (110K) | 18.016.4 | 40.00.0 | 0.00.0 | 8.72.3 | 20.00.0 |
| mixed (220K) | 18.014.8 | 40.00.0 | 0.00.0 | 10.71.2 | 20.00.0 |
| Model | Hard (Original) | Easy | ||
| Paused | Real-time | Paused | Real-time | |
| gpt-4o-mini (19.6 sec) | 75.00.0 | 0.00.0 | 100.00.0 | 0.00.0 |
| gpt-4o (27.2 sec) | 100.00.0 | 0.00.0 | 100.00.0 | 100.00.0 |
| gemini-2.5-pro (99.2 sec) | 100.00.0 | 0.00.0 | 100.00.0 | 0.00.0 |
5.7 Latency-Aware Evaluation in Real-Time Gameplay
Among the 12 games in Orak, 3 games, including Street Fighter III, Super Mario, and Starcraft II, require pausing for evaluation. Here, we conduct a latency-aware evaluation on StarCraft II by enabling real-time gameplay. As shown in Table 8(b), when switching from the paused setting to the real-time mode, all models failed (0 score) under the hard level (the original setup), indicating that timing-sensitive games are strongly affected by inference latency. When the difficulty level was reduced to easy, all models achieved perfect scores under the paused setting, but only GPT-4o maintained perfect performance under real-time conditions. The average response latencies per step with the reflection–planning agent were 19.6 s for GPT-4o-mini, 27.2 s for GPT-4o, and 99.2 s for Gemini-2.5-pro. These results suggest that GPT-4o provides the best trade-off between latency and gameplay accuracy among the three evaluated models.
@@ Conclusion.
We note that Orak provides environments favorable for LLM reasoning and decision-making by pre-processing game states into structured text where information unnecessary to gameplay is hidden. This may offer insights for providers seeking to deploy resource-efficient LLMs on games. We leave a complementary direction, providing the full in-game states with rich uncurated texts to LLMs, for future work. Further discussions, including cost and licensing, are provided in Appendix U.
@@ C.1 Game Description for Street Fighter III
Observation-to-Text Conversion. The Diambra environment offers a convenient interface for extracting the game state from Street Fighter III. Through this interface, we obtain the latest game frame at a resolution of 224×384, along with key state information such as remaining time, player and opponent health, super bar gauge, super count, stun bar gauge, and stun status. However, a critical aspect in fighting games is understanding the relative positions of the characters, which is not directly provided by the Diambra environment. To address this limitation, we employ a lightweight YOLOv11 object detection model (Khanam and Hussain, 2024) to extract the relative positions of the two characters from the game frame. Based on the computed distance, we classify the spatial relationship into three discrete categories—very close, close, and far—and incorporate this information into the agent’s user prompt.
@@ C.3 Evaluation Metric for Street Fighter III
We adopt a Bradley-Terry model formulation (Bradley and Terry, 1952) for Elo estimation, where each model’s rating is iteratively optimized using gradient ascent on the log-likelihood of observed outcomes. The gradient is computed based on the expected win probabilities derived from current ratings, using the standard Elo transformation:
@@ C.5 Result for Street Fighter III
All reported results in Tables 12–13 are computed as the mean and standard deviation over five independent runs, using a ‘zero-shot’ agent for each model configuration. To establish a baseline for comparison, we additionally included a random agent that selects an arbitrary action uniformly at random. This agent was evaluated over 30 episodes, and its performance is summarized in tables.
@@ I.5 Result for Minecraft
Using these difference-based scores, we then compute Elo ratings for all agents following the Bradley-Terry model (Bradley and Terry, 1952), following the approach used in StreetFighter III multi-agent evaluation in Section C.3.
@@
All reported results in Tables 46–46 are computed as the mean and standard deviation over five independent runs, using a ‘zero-shot’ agent for each model configuration. To establish a baseline for comparison, we additionally included a random agent that selects one of the four possible actions (up, down, left, right) uniformly at random. This agent was evaluated over 50 episodes, and its performance is summarized in Tables.